What is the critical angle for light traveling from diamond into water?



1
Expert's answer
2020-12-31T08:08:31-0500

By definition, the critical angle is:


"\theta_c = \arcsin\left(\dfrac{n_2}{n_1} \right)"

where "n_1 = 2.417" is the refractive index of diamond, and "n_2 = 1.333" is the refractive index of water. Thus, obtain:


"\theta_c = \arcsin\left(\dfrac{1.333}{2.417} \right) \approx 33.5\degree"

Answer. "33.5\degree".